Ben earns by giving guitar lessons at 20 per hour. If he spends 10 hours planting ornamental plants and buys 100 worth of pots and cuttings, ehat are the opportunity costs? Compute the accounting profit and economic profit if Ben sold the plants at 200.

1
Expert's answer
2021-12-06T16:40:55-0500

The opportunity costs are 10×20 = 200.

The accounting profit is AP = 200 - 100 = 100.

And economic profit is:

EP = 100 - 200 = -100.